# Topology-Informed Graph Transformer

### v1

@@ Abstract
Transformers have revolutionized performance in Natural Language Processing and Vision, paving the way for their integration with Graph Neural Networks (GNNs). One key challenge in enhancing graph transformers is strengthening the discriminative power of distinguishing isomorphisms of graphs, which plays a crucial role in boosting their predictive performances. To address this challenge, we introduce ’Topology-Informed Graph Transformer (TIGT)’, a novel transformer enhancing both discriminative power in detecting graph isomorphisms and the overall performance of Graph Transformers. TIGT consists of four components: A topological positional embedding layer using non-isomorphic universal covers based on cyclic subgraphs of graphs to ensure unique graph representation: A dual-path message-passing layer to explicitly encode topological characteristics throughout the encoder layers: A global attention mechanism: And a graph information layer to recalibrate channel-wise graph features for better feature representation. TIGT outperforms previous Graph Transformers in classifying synthetic dataset aimed at distinguishing isomorphism classes of graphs. Additionally, mathematical analysis and empirical evaluations highlight our model’s competitive edge over state-of-the-art Graph Transformers across various benchmark datasets.
@@ 1 INTRODUCTION
Transformers have achieved remarkable success in domains such as Natural Language Processing (Vaswani et al., 2023) and Computer Vision (Dosovitskiy et al., 2021). Motivated by their prowess, researchers have applied them to the field of Graph Neural Networks (GNNs). They aimed to surmount the limitations of Message-Passing Neural Networks (MPNNs), which are a subset of GNNs, facing challenges such as over-smoothing (Oono & Suzuki, 2021), over-squashing (Alon & Yahav, 2021), and restricted expressive power (Xu et al., 2019; Morris et al., 2021). An exemplary application of the integration of Transformers into the GNNs field is the Graph Transformer. Multi-head attention mechanism of Transformers is applied to each node in the graph treating the entire set of nodes as if they are fully connected or treating the set of nodes if they are connected by edges. These approaches often come with a low inductive bias, making them prone to over-fitting. Consequently, several implementations blend Graph Transformers with or without MPNNs, yielding promising outcomes (Yang et al., 2021; Ying et al., 2021; Dwivedi & Bresson, 2021; Chen et al., 2022; Hussain et al., 2022; Zhang et al., 2023; Ma et al., 2023; Rampášek et al., 2023; Kong et al., 2023; Zhang et al., 2022).
@@
While Graph Transformers have marked considerable advancements, enhancing Graph Transformers through strengthening the discriminative power of distinguishing isomorphisms of graphs is remaining challenge, which plays a crucial role in boosting their graph-level predictive performances. Previous research has explored various techniques to address the limitations of discriminative power. For instance, studies based on MPNN have enhanced node attributes using high-dimensional complexes, persistent homological techniques, and recurring subgraph structures (Carrière et al., 2020; Bodnar et al., 2021b; Bouritsas et al., 2021; Wijesinghe & Wang, 2021; Bevilacqua et al., 2022; Park et al., 2022; Horn et al., 2022; Choi et al., 2023). Similarly, recent research on Graph Transformers has investigated the use of positional encoding grounded in random walk strategies, Laplacian PE, node degree centrality, and shortest path distance to address these limitations. Furthermore, structure encoding based on substructure similarity has been introduced to amplify the inductive biases inherent in the Transformer.
This paper introduces a Topology-Informed Graph Transformer(TIGT), which embeds sophisticated topological inductive biases to augment the model’s expressive power and predictive efficacy. Before the Transformer layer, each node attribute is integrated with a topological positional embedding layer based on the differences of universal covers obtained from the original graph structure and collections of unions of cyclic subgraphs, the topological invariants of which contains their first homological invariants.
@@
In companion with the novel positional embedding layer, we explicitly encode cyclic subgraphs in the dual-path message passing layer and incorporate channel-wise graph information in the graph information layer. These are combined with global attention across all Graph Transformer layers, drawing inspiration from (Choi et al., 2023) and (Rampášek et al., 2023). As a result, the TIGT layer can concatenate hidden representations from the dual-path message passing layer, combining information of original structure and cyclic subgraphs, global attention layer, and graph information layer to preserve both topological information and graph-level information in each layer. Specifically, the dual-path message passing layer enables overcoming the limitations of positional encoding and structural encoding to increase expressive power when the number of layers increases. We justify the proposed model’s expressive power based on the theory of covering space. Furthermore, we perform experiment in synthetic datasets aimed at distinguishing isomorphism classes of graphs and benchmark datasets aimed at demonstrating the state-of-the-art of competitive predictive performance of the proposed model.
@@
Our main contributions can be summarized as follows: (i) Theoretical justification of expressive powers of TIGT and its comparison with other Graph Transformers by utilizing the theory of covering spaces, comparison of Euler characteristic formulae of graphs and their subgraphs, and the geometric rate of convergence of Markov operators over finite graphs to stationary distributions. (ii) Novel positional embedding layer based on the MPNNs and simple architectures to enrich topological information in each Graph Transformer layer (iii) Outperformance shown in processing synthetic dataset to assess the expressive power of GNNs (iv) State-of-art or competitive results, especially in the large graph-level benchmarks.
@@ 2 Preliminary
Message passing neural networks MPNNs have demonstrated proficiency in acquiring vector representations of graphs by handling local information based on the connectivity between nodes among other types of GNNs such as Graph Convolution Network (GCN), Graph Attention Network (GAT) (Veličković et al., 2018), Graph Isomorphism Network (GIN) (Xu et al., 2019) and Residual Graph ConvNets (GatedGCN) (Bresson & Laurent, 2018). We denote $\text{MPNN}^{l}$ when it has a composition of $l$ neighborhood aggregatin layers. Each $l$-th layer $H^{(l)}$ of the network constructs hidden node attributes of dimension $k_{l}$, denoted as $h_{v}^{(l)}$, using the following composition of functions:
@@
where $X_{v}$ is the initial node attribute at $v$. Let $M_{v}^{(l)}$ be the collection of all multisets of $k_{l-1}$-dimensional real vectors with $\deg v$ elements counting multiplicities.
@@
is a set theoretic function of $k_{l}^{\prime}$-dimensional real vectors, and the combination function
@@
Clique adjacency matrix The clique adjacency matrix, proposed by  (Choi et al., 2023), is a matrix that represents bases of cycles in a graph in a form analogous to the adjacency matrix, enabling its processing within GNNs. Extracting bases of cycles results in incorporating a topological property equivalent to the first homological invariants of graphs (Paton, 1969). The set of cyclic subgraphs of $G$ which forms the basis of the cycle space (or the first homology group) of $G$ is defined as the cycle basis $\mathcal{B}_{G}$. The clique adjacency matrix, $A_{C}$, is the adjacency matrix of the union of #$\#\mathcal{B}_{G}$ complete subgraphs, each obtained from adding all possible edges among the set of nodes of each basis element $B\in\mathcal{B}_{G}$. Explicitly, the matrix $A_{C}:=\{a^{C}_{u,v}\}_{u,v\in V(G)}$ is given by
@@ 3 TOPOLOGY-INFORMED GRAPH TRANSFORMER(TIGT)
In this section, we introduce overall TIGT architecture. The overall architecture of our model is illustrated in Figure 1.Suppose we are given the graph $G:=(V,E)$. It can be represented by four types of matrices to use input of TIGT; a node feature matrix $X\in\mathbb{R}^{n\times k_{X}}$, an adjacency matrix $A\in\mathbb{R}^{n\times n}$ , a clique adjacency matrix $A_{c}\in\mathbb{R}^{n\times n}$ and an edge feature matrix $E\in\mathbb{R}^{n\times k_{E}}$. Note that $n$ is the number of node, $k_{X}$ is node feature dimension and $k_{E}$ is edge feature dimension. The clique adjacency matrices are obtained from same process in previous research (Choi et al., 2023). For clarity and conciseness in our presentation, we have omitted the details pertaining to the normalization layer and the residual connection. We note that some of the mathematical notations used in explaining the model design and details of model structures conform to those shown in (Rampášek et al., 2023).
@@ 3.1 Topological positional embedding layer
Most of previous researches of Graph Transformers use positional embeddings based on parameters such as node distance, random walk, or structural similarities, in order to adequately capture the nuances of graph structures. Diverging from this typical approach, we propose a novel method for obtaining learnable positional encodings by leveraging MPNNs. This approach aims to enhance the discriminative power with respect to isomorphism classes of graphs, drawing inspiration from Cy2C-GNNs (Choi et al., 2023). First, we use any MPNNs to obtained hidden attribute from original graph and new graph structure with clique adjacency matrix as follows:
@@
where $X$ represents the node embedding tensor from the embedding layers, and [][\ \ ][ ] denotes the process of stacking two hidden representations. It’s important to note that MPNN for adjacency matrix and clique adjacency matrix share weights, and the number of layers of MPNNs are not a constraint. Then the node features are updated along with topological positional attributes, as shown below:
@@
where $i$ is the node index in graph and $\theta_{pe}\in R^{1\times k\times 2}$ represents the learnable parameters that are utilized to integrate features from two different universal cover. The SUM operation performs a sum of the hidden features $h_{A}$ and $h_{A_{c}}$ by summing over the last dimensions. For the Activation function, in this study, we use hyperbolic tangent function to bound the value of positional information. Regardless of the presence or absence of edge attributes, we do not use any existing edge attributes in this layer. The main objective of this layer is to enrich node features by adding topological information by combining the two universal covers. The results $X^{0}$ will be subsequently fed into the encoder layers of TIGT.
@@ 3.2 Encoder layer of TIGT
The Encoder layer of the TIGT is based on the three components: A Dual-path message passing layer: A global attention layer: And a graph information layer. For the input to the Encoder layer, the transformed input feature $X^{l-1}$, along with $A$, $A_{c}$, and $E^{l-1}$ is utilized, where $l$ is the encoder layer number in TIGT.
Dual-path MPNNs     Hidden representations $X^{l-1}$, sourced from the preceding layer, paired with the adjacency matrix and clique adjacency matrix, are processed through a dual-path message passing layer as follows:
@@
Global attention layer To capture global relationship of each node, we apply multi-head attention of vanilla Transformer as follows:
@@
where MHA is multi-head attention layer. Then we obtained representation vectors from local neighborhood, all nodes in graph and neighborhood in same cyclic subgraph. Combining these representations, we obtain the intermediate node representations given by:
@@
Graph information layer    The pooled graph features, extracted from the Graph Information Layer, are seamlessly integrated. Inspired by the squeeze-and-excitation block (Hu et al., 2019), this process adaptively recalibrates channel-wise graph features into each node feature as:
@@ 3.3 Mathematical background of TIGT
Clique adjacency matrix     The motivation for utilizing the clique adjacency matrix in implementing TIGT originates from the recent work by Choi et al., which establishes a mathematical identification of discerning capabilities of GNNs using the theory of covering spaces of graphs (Choi et al., 2023). To summarize their work, conventional GNNs represent two graphs $G$ and $H$, endowed with node feature functions $X_{G}:G\to\mathbb{R}^{k}$ and $X_{H}:G\to\mathbb{R}^{k}$, as identical vector representations if and only if the universal covers of $G$ and $H$ are isomorphic, and the pullback of node attributes over the universal covers are identical. We note that universal covers of graphs are infinite graphs containing unfolding trees of the graph rooted at a node as subgraphs. In other words, these universal covers do not contain cyclic subgraphs which may be found in the original given graph as subgraphs. Additional measures to further distinguish cyclic structures of graphs are hence required to boost the distinguishing power of GNNs. Among various techniques to represent cyclic subgraphs, we focus on the following two solutions which can be easily implemented: (1) Construct clique adjacency matrices $A_{C}$, as utilized in the architectural component of TIGT, which transform the geometry of universal covers themselves: (2) Impose additional positional encodings, which alter the pullback of node attributes over the universal covers. The distinguishing power of TIGT can be stated as follows, whose proof follows from the results shown in (Choi et al., 2023).
@@
Suppose $G$ and $H$ are two graphs with the same number of nodes and edges. Suppose that there exists a cyclic subgraph $C$ that is an element of a cycle basis of $G$ such that satisfies the following two conditions: (1) $C$ does not contain any proper cyclic subgraphs (2) Any element of a cycle basis of $H$ is not isomorphic to $C$. Then TIGT can distinguish $G$ and $H$ as non-isomorphic.
@@
There exists a pair of graphs $G$ and $H$ such that TIGT can distinguish as non-isomorphic whereas 3-Weisfeiler-Lehman (3-WL) test cannot.
@@
Graph biconnectivity     Given that TIGT is able to distinguish classes of graphs that 3-WL cannot, it is reasonable to ask whether TIGT can capture topological properties of graphs that state-of-the-art techniques can encapsulate, which is the problem of detecting bi-connectivity of graphs (Zhang et al., 2023; Ma et al., 2023). We recall that a connected graph $G$ is vertex (or edge) biconnected if there exists a vertex $v$ (or an edge $e$) such that $G\setminus\{v\}$ (or $G\setminus\{e\}$) has more connected components than $G$. As these state-of-the-art techniques can demonstrate, TIGT as well is capable to distinguish vertex (or edge) bi-connectivity. The idea of the proof relies on comparing the Euler characteric formula for graphs $G$ and $G\setminus\{v\}$ (or $G\setminus\{e\})$, the specific details of which are provided in Appendix A.3.
@@
In relation to the above observation, we show that the relative random walk probabilities positional encoding (RRWP) suggested in (Ma et al., 2023) may not effectively model $K$ steps of random walks on graphs $G$ containing a cyclic subgraph with odd number of nodes and may not be distinguishable by $1$-WL as $K$ grows arbitrarily large, the proof of which is outlined in Appendix A.4.
@@
Let $\mathcal{G}$ be any collections of graphs whose elements satisfy the following three conditions: (1) All graphs $G\in\mathcal{G}$ share the same number of nodes and edges: (2) Any $G\in\mathcal{G}$ contains a cyclic subgraph with odd number of nodes: (3) For any number $d\geq 1$, all the graphs $G\in\mathcal{G}$ have identical number of nodes whose degree is equal to $d$. Fix an integer $K$, and suppose the node indices for $G\in\mathcal{G}$ are ordered based on its increasing degrees. Let $\mathbf{P}$ be the RRWP positional encoding associated to $G$ defined as $\mathbf{P}_{i,j}:=[\textbf{I},\textbf{M},\textbf{M}^{2},\cdots,\textbf{M}^{K-1}]_{i,j}\in\mathbb{R}^{K}$, where $\textbf{M}:=\textbf{D}^{-1}\textbf{A}$ with A being the adjacency matrix of $G$, and D the diagonal matrix comprised of node degrees of $G$. Then there exists a unique vector $\pi\in\mathbb{R}^{n}$ independent of the choice of elements in $\mathcal{G}$ and a number $10<\gamma<$ 1 such that for any $10\leq l\leq K-$ 1, we have $\max_{(i,j)}\|\textbf{M}^{l}_{i,j}-\pi_{j}\|<\gamma^{l}$.
In particular, the theorem states that the positional encodings which are intended to model $K$ steps of random walks converge at a geometric rate to a fixed encoding $\pi\in\mathbb{R}^{K}$ regardless of the choice of non-isomorphism classes of graphs $G\in\mathcal{G}$. Hence, such choices of positional encodings may not be effective enough to represent differences in topoloical structures among such graphs as differences in their vector representations.

### v2

@@ Abstract
Transformers, through their self-attention mechanisms, have revolutionized performance in Natural Language Processing and Vision. Recently, there has been increasing interest in integrating Transformers with Graph Neural Networks (GNNs) to enhance analyzing geometric properties of graphs by employing global attention mechanisms. A key challenge in improving graph transformers is enhancing their ability to distinguish between isomorphic graphs, which can potentially boost their predictive performance. To address this challenge, we introduce ’Topology-Informed Graph Transformer (TIGT)’, a novel transformer enhancing both discriminative power in detecting graph isomorphisms and the overall performance of Graph Transformers. TIGT consists of four components: (1) a topological positional embedding layer using non-isomorphic universal covers based on cyclic subgraphs of graphs to ensure unique graph representation, (2) a dual-path message-passing layer to explicitly encode topological characteristics throughout the encoder layers, (3) a global attention mechanism, and (4) a graph information layer to recalibrate channel-wise graph features for improved feature representation. TIGT outperforms previous Graph Transformers in classifying synthetic dataset aimed at distinguishing isomorphism classes of graphs. Additionally, mathematical analysis and empirical evaluations highlight our model’s competitive edge over state-of-the-art Graph Transformers across various benchmark datasets.
@@ 1 INTRODUCTION
Transformers have achieved remarkable success in domains such as Natural Language Processing (Vaswani et al., 2023) and Computer Vision (Dosovitskiy et al., 2021). Motivated by their prowess, researchers have applied them to the field of Graph Neural Networks (GNNs). They aimed to surmount the limitations of Message-Passing Neural Networks (MPNNs), a subset of GNNs, by attempting to address challenges such as over-smoothing (Oono & Suzuki, 2021), over-squashing (Alon & Yahav, 2021), and restricted expressive power (Xu et al., 2019; Morris et al., 2021). An exemplary application of the integration of Transformers into GNNs is the Graph Transformer. One way to achieve this integration is by applying multi-head attention mechanism of Transformers to each node of a given graph dataset. Such a technique treats the set of all nodes as being fully interconnected or connected by edges. This approach, however, often has limited capabilities in guaranteeing a strong inductive bias  [^1] For example, unless trained over large scale data, transformers may lack inherent inductive biases in modeling permutation and translational invariance, hierarchical structures, and local data structures  (Hahn, 2020; Dosovitskiy et al., 2021; Xu et al., 2021), hence making it prone to overfitting. To address this drawback, previous studies focused on devising new implementations to blend Graph Transformers with other deep learning techniques including MPNNs, thereby yielding promising empirical results  (Yang et al., 2021; Ying et al., 2021; Dwivedi & Bresson, 2021; Chen et al., 2022; Hussain et al., 2022; Zhang et al., 2023; Ma et al., 2023; Rampášek et al., 2023; Kong et al., 2023; Zhang et al., 2022).
@@
While Graph Transformers boast considerable advancements, enhancing them by strengthening their discriminative powers in distinguishing isomorphism classes of graphs still remains a challenge, an approach that can potentially boost their performance in predicting various properties of graph datasets. For instance, studies based on MPNN have enhanced node attributes by utilizing high-dimensional complexes, persistent homological techniques, and recurring subgraph structures (Carrière et al., 2020; Bodnar et al., 2021b; Bouritsas et al., 2021; Wijesinghe & Wang, 2021; Bevilacqua et al., 2022; Park et al., 2022; Horn et al., 2022; Choi et al., 2023; Li et al., 2023; Choi et al., 2024; Zhou et al., 2024). Other approaches addressing these limitations include incorporating positional encoding grounded in random walk strategies, Laplacian Positional Encoding (PE), node degree centrality, and shortest path distance. Furthermore, structure encoding based on substructure similarity has been introduced to amplify the inductive biases inherent in the Transformer.
@@
This paper introduces a Topology-Informed Graph Transformer (TIGT), which embeds topological inductive biases to augment the model’s expressive power and predictive efficacy. Prior to the Transformer layer, TIGT augments each node attribute with a topological positional embedding layer based on the differences of universal covers (or unfolding trees) obtained from the original graph and collections of unions of its cyclic subgraphs. These new topological biases allow TIGT to encapsulate the first homological invariants (or cyclic structures) [^2] Given a graph $G=(V,E)$, the first homology group of $G$ is a free abelian group on a cycle basis of $G$. A cycle basis of $G$ is a minimal set of cyclic subgraphs of $G$ such that all cyclic subgraphs of $G$ can be expressed as unions or complements of elements in the cycle basis. For a more rigorous treatment on the construction of homology groups of topological spaces in general, we refer to Chapter 2 of  (Hatcher, 2002). of the given graph datasets. In combination with the novel positional embedding layer, we explicitly encode cyclic subgraphs in the dual-path message passing layer and incorporate channel-wise graph information in the graph information layer. These are combined with global attention across all Graph Transformer layers, a design of which was inspired from (Choi et al., 2023) and (Rampášek et al., 2023). As a result, the TIGT layer can concatenate hidden representations from the dual-path message passing layer, thereby combining information of the original graph dataset and its cyclic subgraphs, global attention layer, and graph information layer. This allows TIGT to preserve both topological information and graph-level information in each layer. Specifically, the dual-path message passing layer enables TIGT to overcome the limitations of positional encoding and structural encoding to increase expressive power when the number of layers increases. We justify the proposed model’s expressive power based on the theory of covering spaces. Furthermore, we perform experiments on synthetic datasets to distinguish isomorphism classes of graphs. We also evaluate the proposed model on benchmark datasets to demonstrate its state-of-the-art predictive performance.
@@
Our main contributions can be summarized as follows. (i). We provide a theoretical justification for the expressive power of TIGT. We compare TIGT with other Graph Transformers using a number of results from geometry and probability theory. These tools include the theory of covering spaces (Hatcher, 2002, Chapter 1.3), Euler characteristic formulas of graphs and their subgraphs (Hatcher, 2002, Chapter 2.2), and the geometric rate of convergence of Markov operators over finite graphs to stationary distributions (Lovasz, 1993). (ii) We propose a novel positional embedding layer based on the MPNNs and simple architectures to enrich topological information in each Graph Transformer layer. (iii). We demonstrate superior performance in processing synthetic datasets to assess the expressive power of GNNs. (iv) We obtain state-of-the-art or competitive results, especially in large graph-level benchmarks.
@@ 2 Preliminary
We first introduce a number of background materials which could be helpful for understanding the newly proposed architecture.
Message passing neural networks. MPNNs have demonstrated proficiency in generating vector representations of graphs by exploiting local information based on node connectivity. This capability is shared with other types of GNNs, such as Graph Convolutional Network (GCN), Graph Attention Network (GAT) (Veličković et al., 2018), Graph Isomorphism Network (GIN) (Xu et al., 2019) and Residual Graph ConvNets (GatedGCN) (Bresson & Laurent, 2018). We use the abbreviation $\text{MPNN}^{l}$ to denote an MPNN that has a composition of $l$ neighborhood aggregating layers. Each $l$-th layer $H^{(l)}$ of the network constructs hidden node attributes of dimension $k_{l}$, denoted as $h_{v}^{(l)}$, using the following composition of functions:
@@
where $X_{v}$ is the initial node attribute at $v$. Let $M_{v}^{(l)}$ be the collection of all multisets of $k_{l-1}$-dimensional real vectors with $\deg v$ elements counting multiplicities. The aggregation function
@@
is a set theoretic function that outputs $k_{l}^{\prime}$-dimensional real vectors, and the combination function
@@
Clique adjacency matrix The clique adjacency matrix, proposed by  (Choi et al., 2023), is a matrix that represents bases of cycles of a graph in a form analogous to the adjacency matrix, enabling its processing within GNNs. Extracting bases of cycles as in  (Paton, 1969), removing edges of the original graph not contained in the cycle bases, and substituting cyclic subgraphs in the cycle bases with cliques result in incorporating topological properties equivalent to first homological invariants (or cyclic substructures) of graphs. The set of cyclic subgraphs of $G$ which forms the basis of the cycle space (or the first homology group) of $G$ is defined as the cycle basis $\mathcal{B}_{G}$. The clique adjacency matrix, $A_{C}$, is the adjacency matrix of the union of #$\#\mathcal{B}_{G}$ complete subgraphs, each obtained from adding all possible edges among the set of nodes of each basis element $B\in\mathcal{B}_{G}$. Explicitly, the matrix $A_{C}:=\{a^{C}_{u,v}\}_{u,v\in V(G)}$ is given by
@@ 3 TOPOLOGY-INFORMED GRAPH TRANSFORMER(TIGT)
In this section, we introduce the overall TIGT architecture. The overall architecture of our model is illustrated in Figure 1. Suppose we are given the graph $G:=(V_{G},E_{G})$, where $V_{G}$ is the set of nodes and $E_{G}$ is the set of edges of $G$. The graph $G$ can be represented by four types of matrices which are used as inputs of TIGT: (1) a node feature matrix $X\in\mathbb{R}^{n\times k_{X}}$, (2) an adjacency matrix $A\in\mathbb{R}^{n\times n}$, (3) a clique adjacency matrix $A_{c}\in\mathbb{R}^{n\times n}$, and (4) an edge feature matrix $E\in\mathbb{R}^{n\times k_{E}}$. Note that $n$ is the number of nodes, $k_{X}$ is node feature dimension, and $k_{E}$ is edge feature dimension. The clique adjacency matrices are obtained using the same procedure outlined in previous research (Choi et al., 2023). For clarity and conciseness in our presentation, we leave the details pertaining to the normalization layer and the residual connection in Appendix C.2. We note that some of the mathematical notations used in explaining the model design and details of model structures conform to those shown in (Rampášek et al., 2023).
@@ 3.1 Topological positional embedding layer
To adequately capture the nuances of graph structures, most previous work in Graph Transformers uses positional embeddings based on parameters such as node distance, random walk, or structural similarities. Diverging from this typical approach, we propose a novel method for obtaining learnable positional encodings by leveraging MPNNs. This approach aims to enhance their discriminative power with respect to isomorphism classes of graphs, drawing inspiration from Cy2C-GNNs (Choi et al., 2023). First, we use a MPNNs to obtain hidden attributes from the original graph and a new graph structure with clique adjacency matrix as follows:
@@
where $X$ represents the node embedding tensor from the embedding layers, and [][\ \ ][ ] denotes the process of stacking two hidden representations. It’s important to note that MPNNs for the adjacency matrix and the clique adjacency matrix share weights. Then the node features are updated along with topological positional attributes, as shown below:
@@
where $i$ is the node index of the original graph and $\theta_{pe}\in R^{1\times k\times 2}$ represents the learnable parameters that are utilized to integrate features from two universal covers (or unfolding trees) of the two graph structures. The SUM operation performs a sum of the hidden features $h_{A}$ and $h_{A_{c}}$ by summing over the last dimensions. For the Activation function, in this study, we use hyperbolic tangent function to bound the value of positional information. Regardless of the presence or absence of edge attributes, we do not use any existing edge attributes in this layer. The main objective of this layer is to enrich node features by adding topological information obtained from combining the two universal covers. The resulting output $X^{0}$ will be subsequently fed into the encoder layers of TIGT.
@@ 3.2 Encoder layer of TIGT
The Encoder layer of the TIGT is based on three components: A Dual-path message passing layer, a global attention layer, and a graph information layer. For the input to the Encoder layer, we utilize the transformed input feature $X^{l-1}$ along with $A$, $A_{c}$, and $E^{l-1}$, where $l$ is encoder layer number in TIGT.
Dual-path MPNNs     Hidden representations $X^{l-1}$, sourced from the preceding layer, are paired with the adjacency matrix and clique adjacency matrix. The paired inputs are then processed through a dual-path message passing layer as follows:
@@
Global attention layer To capture global relationship among nodes of the original graph, we apply the multi-head attention of the vanilla Transformer as follows:
@@
where MHA is multi-head attention layer. Then we obtain representation vectors from local neighborhoods, all nodes in graph, and neighborhoods of the same cyclic subgraph. Combining these representations, we obtain the intermediate node representations given by:
@@
Graph information layer     We propose a Graph Information Layer to integrate the pooled graph features with the output of the global attention layer. Inspired by the squeeze-and-excitation block (Hu et al., 2019), this process adaptively recalibrates channel-wise graph features into each node feature as:
@@ 3.3 Mathematical background of TIGT
Now that we have introduced the architectural design of TIGT, we focus on elucidating mathematical insights which enable TIGT to exhibit competitive discriminative power in comparison to other state-of-the-art techniques.
Clique adjacency matrix     The motivation for utilizing the clique adjacency matrix in implementing TIGT originates from recent work by previous research (Choi et al., 2023), which establishes a mathematical identification of discerning capabilities of GNNs using the theory of covering spaces of graphs. To summarize their work, conventional GNNs represent two graphs $G$ and $H$, endowed with node feature functions $X_{G}:G\to\mathbb{R}^{k}$ and $X_{H}:G\to\mathbb{R}^{k}$, as identical vector representations if and only if the universal covers of $G$ and $H$ are isomorphic, and the pullback of node attributes over the universal covers are identical. We note that universal covers of graphs are infinite graphs containing unfolding trees of the graph rooted at a node as subgraphs. In other words, these universal covers do not contain cyclic subgraphs which may be found in the original given graph as subgraphs. Additional measures to further distinguish cyclic structures of graphs are hence required to boost the distinguishing power of GNNs. Among various techniques to represent cyclic subgraphs, we focus on the following two solutions which can be easily implemented. (1) Construct clique adjacency matrices $A_{C}$, as utilized in the architectural component of TIGT, which transform the geometry of universal covers themselves. (2) Impose additional positional encodings, which alter the pullback of node attributes over the universal covers. The distinguishing power of TIGT can be stated as follows, whose proof follows from the results shown in (Choi et al., 2023).
@@
Suppose $G$ and $H$ are two graphs with the same number of nodes and edges. Suppose that there exists a cyclic subgraph $C$ that is an element of a cycle basis of $G$ such that satisfies the following two conditions: (1) $C$ does not contain any proper cyclic subgraphs, and (2) any element of a cycle basis of $H$ is not isomorphic to $C$. Then TIGT can distinguish $G$ and $H$ as non-isomorphic.
@@
There exists a pair of graphs $G$ and $H$ such that TIGT can distinguish them to be non-isomorphic, whereas 3-Weisfeiler-Lehman (3-WL) test cannot.
@@
Graph biconnectivity     Given that TIGT is able to distinguish classes of graphs that 3-WL cannot, it is reasonable to ask whether TIGT can capture topological properties of graphs that other state-of-the-art techniques can encapsulate. One of such properties is bi-connectivity of graphs (Zhang et al., 2023; Ma et al., 2023). We recall that a connected graph $G$ is vertex (or edge) biconnected if there exists a vertex $v$ (or an edge $e$) such that $G\setminus\{v\}$ (or $G\setminus\{e\}$) has more connected components than $G$. As these state-of-the-art techniques can demonstrate, TIGT as well is capable to distinguish vertex (or edge) bi-connectivity of graphs. The idea of the proof relies on comparing the Euler characteric formula for graphs $G$ and $G\setminus\{v\}$ (or $G\setminus\{e\})$, the specific details of which are provided in Appendix A.3.
@@
In relation to the above observation, we show that the relative random walk probabilities positional encoding (RRWP) suggested in (Ma et al., 2023) may not effectively model $K$ steps of random walks on graphs $G$ containing a cyclic subgraph with odd number of nodes, and may not be distinguishable by $1$-WL as $K$ grows arbitrarily large. The proof of the theorem is in Appendix A.4.
@@
Let $\mathcal{G}$ be any collections of graphs whose elements satisfy the following three conditions: (1) All graphs $G\in\mathcal{G}$ share the same number of nodes and edges, (2) any $G\in\mathcal{G}$ contains a cyclic subgraph with odd number of nodes, and (3) for any number $d\geq 1$, all the graphs $G\in\mathcal{G}$ have identical number of nodes whose degree is equal to $d$. Fix an integer $K$, and suppose the node indices for $G\in\mathcal{G}$ are ordered based on its increasing degrees. Let $\mathbf{P}$ be the RRWP positional encoding associated to $G$ defined as $\mathbf{P}_{i,j}:=[\textbf{I},\textbf{M},\textbf{M}^{2},\cdots,\textbf{M}^{K-1}]_{i,j}\in\mathbb{R}^{K}$, where $\textbf{M}:=\textbf{D}^{-1}\textbf{A}$ with A being the adjacency matrix of $G$, and D the diagonal matrix comprised of node degrees of $G$. Then there exists a unique vector $\pi\in\mathbb{R}^{n}$ independent of the choice of elements in $\mathcal{G}$ and a number $10<\gamma<$ 1 such that for any $10\leq l\leq K-$ 1, we have $\max_{(i,j)}\|\textbf{M}^{l}_{i,j}-\pi_{j}\|<\gamma^{l}$.
In particular, the theorem states that the positional encodings which are intended to model $K$ steps of random walks converge at a geometric rate to a fixed encoding $\pi\in\mathbb{R}^{K}$ regardless of the choice of non-isomorphism classes of graphs $G\in\mathcal{G}$. Hence, such choices of positional encodings may not be effective enough to represent differences in topological structures among such graphs as differences in their vector representations.